OpenStudy (anonymous):

For what value of θ, in degrees, is sin θ = cos 58°?

OpenStudy (anonymous):

Cofunction identities sin (90° - Θ) = cos Θ cos (90° - Θ) = sin Θ Θ = 90° - 58°

OpenStudy (anonymous):

Or you could take the arcsin. \[\theta = \sin^{-1} (\cos 58°)\]